OpenStudy (anonymous):

find dy/dx of y=1/x using (f(x+h) -f(x)) / h. I need step by step explaination. I keep getting really lost in all the fractions

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

you really want to understand this and do it step by step? we can if you like

OpenStudy (anonymous):

i havent got an answer with explaination yet. i know the answer i just cant do the formula.

OpenStudy (anonymous):

ok we can work it out slowly. lets start with \[\frac{1}{3}-\frac{1}{2}\]

OpenStudy (anonymous):

because the only part that requires any work is subtracting the fractions. i am not trying to insult your intelligence, it is just that if we do it for numbers we will see why it works here

OpenStudy (anonymous):

or would you prefer if i just write the steps necessary?

OpenStudy (anonymous):

Its ((1/x+h)-(1/x))/h but then what do you multiply top and bottom by?

OpenStudy (anonymous):

nothing. we will actually subtract in the numerator

OpenStudy (anonymous):

\[\frac{1}{a}-\frac{1}{b}=\frac{b-a}{ab}\]

OpenStudy (anonymous):

so \[\frac{1}{x+h}-\frac{1}{x}=\frac{x-(x+h)}{(x+h)x}\]

OpenStudy (anonymous):

therefore the numerator of the "difference quotient" is \[\frac{-h}{(x+h)x}\]

OpenStudy (anonymous):

now we want to divide this by h right? and when we divide by h we get \[\frac{-1}{(x+h)x}\]

OpenStudy (anonymous):

now you take the limit as h goes to zero by simply replacing h by 0 in the above expression and get \[\frac{-1}{(x+0)x}=-\frac{1}{x^2}\]

OpenStudy (anonymous):

any step not clear?

OpenStudy (anonymous):

I've never heard of the b-a/ab. is that just something your supposed to know? is there no other way to do it?

OpenStudy (anonymous):

that is how i subtract fractions. more genally it is \[\frac{a}{b}-\frac{c}{d}=\frac{ad-bc}{bd}\] but in this case the numerators are 1. it is not some special formula, it is just how one subtracts fractions

OpenStudy (anonymous):

that is why i wanted to start with \[\frac{1}{3}-\frac{1}{2}\] so we remember it is \[\frac{2-3}{3\times 2}=\frac{-1}{6}\]

OpenStudy (anonymous):

likewise \[\frac{1}{5}-\frac{1}{4}=\frac{-1}{20}\]

OpenStudy (anonymous):

what happened to the orignial h in the denomenator?

OpenStudy (anonymous):

once we got \[\frac{-h}{(x+h)x}\] that was just the numerator of the "difference quotient" we still had to divide by h

OpenStudy (anonymous):

it was \[\frac{f(x+h)-f(x)}{h}=\frac{\frac{-h}{(x+h)x}}{h}\]

OpenStudy (anonymous):

when you divide by h, the h's cancel. that is why you end up with just \[\frac{-1}{(x+h)x}\]

OpenStudy (anonymous):

more clear now?